OpenStudy (anonymous):

cos^4(x) - sin^4(x) = 1 - 2sin^2(x)

OpenStudy (raden):

use the difference square formula to simplifying the letf side. a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

Would this be right then for LS? (cosx - sinx)(cosx + sinx) ?

OpenStudy (raden):

so, cos^4(x) - sin^4(x) = (cos^2(x) - sin^2(x)) (cos^2(x) + sin^2(x) = (cos^2(x) - sin^2(x)) then use the identity ; cos^2 (x) = 1 - sin^2 (x) therefore, (cos^2(x) - sin^2(x)) = 1 - sin^2 (x) - sin^2 (x) = 1 - 2sin^2(x)

OpenStudy (anonymous):

what did you do from = (cos^2(x) - sin^2(x)) (cos^2(x) + sin^2(x) = (cos^2(x) - sin^2(x))

OpenStudy (raden):

this is an identity also, cos^2 x + sin^2 = 1 (cos^2(x) - sin^2(x)) (cos^2(x) + sin^2(x) = (cos^2(x) - sin^2(x)) * 1 = (cos^2(x) - sin^2(x))

OpenStudy (anonymous):

Ah I see ok great!